ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 169, 2009

Androulakis, G., P. Dodos, G. Sirotkin and V. G. Troitsky Classes of strictly singular operators and their products

221

Ben-Ari, I. The asymptotic shift for the principal eigenvalue for second order elliptic operators in the presence of small obstacles

181

Björner A., M. Wachs and V. Welker On sequentially Cohen-Macaulay complexes and posets

295

Calegari, F. and M. Emerton Elliptic curves of odd modular degree

417

Castillo, J. M. F. and J. Suárez Extending operators into Lindenstrauss spaces

1

Cheng, X. and Z. Shen A class of Finsler metrics with isotropic S-curvature

317

Djorić, M. and M. Okumura Certain condition on the second fundamental form of CR submanifolds of maximal CR dimension of complex projective space

47

Dodos, P. See G. Androulakis Emerton, M. See F. Calegari Gatzouras, D. and A. Giannopoulos Threshold for the volume

spanned by random points with independent coordinates

125

Giannopoulos, A. See D. Gatzouras

Guersenzvaig, N. H. Elementary criteria for irreducibility of f(X^{r})  109

Hanzer, M. The unitarizability of the Aubert dual of strongly positive square integrable representations

251

Hjorth, G. and B. D. Miller Ends of graphed equivalence relations, II

393

Li, C. On the distribution of Satake parameters of GL_{2} holomorphic cuspidal representations

341

Lu, D., Q. Wu and J. J. Zhang Hopf algebras with rigid dualizing complexes

89

Miller, B. D. See G. Hjorth

Miller, B. D. Ends of graphed equivalence relations, I  375
Moreno, Y. and A. Plichko On automorphic Banach spaces  29

Okumura, M. See M. Djorić Plichko, A. See Y. Moreno

Regev, A. Probabilities in the (k, ℓ) hook  61

Rivoal, T. and M. Welter Sur les fonctions arithmétiques non entières

155

Shen, Z. See X. Cheng Sirotkin, G. See G. Androulakis Suárez, J. See J. M. F. Castillo Troitsky, V. G. See G. Androulakis Wachs, M. See Björner Anders Welker, V. See Björner Anders Welter, M. See T. Rivoal Wu, Q. See D. Lu Zhang, J. J. See D. Lu